By mass, one alloy consists of 60% copper and 40% ziinc. Another is 30% copper and 70% zinc. How many grams of each alloy are needed to make 120g of an alloy that consists of 50% of each metal?
Answer by stanbon(75887)

By mass, one alloy consists of 60% copper and 40% zinc. Another is 30% copper and 70% zinc. How many grams of each alloy are needed to make 120g of an alloy that consists of 50% of each metal?
--------
Copper Eq:: 0.6x + 0.3y = 0.5*120
Zinc Eq:: 0.4x + 0.7y = 0.5*120
----------------------
6x + 3y = 5*120
4x + 7y = 5*120
------
24x + 12y = 20*120
24x + 42y = 30*120
-----
30y = 1200
y = 40 grams (wt. of alloy with 30% copper)
---
x = 120-40 = 80 grams (wt of alloy with 60% copper)
----------
Cheers,
Stan H.
-----------